Can you please help me with this equation: 6y+7=3+2(y-4)
Answer by checkley71(8403)

6Y+7=3+2(Y-4) OR 6Y+7=3+2Y-8 OR 6Y-2Y=3-8-7 OR 4Y=-12 OR Y=-12/4 OR Y=-3
PROOF 6(-3)+7=3+2(-3-4) OR -18+7=3+2(-7) OR -11=3-14 OR -11=-11